# Investigating the Fermi-Hubbard model by the Tensor-Backflow method

### v2

@@ III Numerical results
Figure 1: Energy extrapolations and comparisons between Tensor-Backflow and fPEPS under OBC for the lattice of $8\times 8$(a) and $4\times 16$(b). Energy extrapolation based on Tensor-Backflow results under PBC for the lattice of $4\times 16$(c). $\sigma^{2}$ is the energy variance. Red and blue solid dots depict Tensor-Backflow with NN and all-site backflow terms. For solid dots of the same color, low and high energies depict results with and without one Lanczos optimization.
III.1 Comparisons with previous results
@@
Fig.(1)(a)(b) depict energy comparisons respect to the fPEPS method under open-boundary condition (OBC) on 64-site lattices. The fPEPS is a state-of-the-art method for latticed fermions and its state representation ability can be improved by increasing the bond-dimension ($D$). From literature Hubbard_fPEPS , the complexity of fPEPS scales rapidly respect to $D$: $\mathcal{O}(D^{6})$, and the maximal bond dimension is $D$=20. The fPEPS is firstly optimized by the imaginary-time-evolution method named as the simple-update (SU), followed by the gradient-optimization (GO). The fPEPS with $D$=20 can achieve extremely high energy precision under OBC.
In our optimizations, the initial state of Tensor-Backflow is unrestricted Hartree-Fock (UHF) state represented by a tensor of the dimension $[M,N,2]$, where $M$ is the site number $N$ is the particle number and 2 is because of spin. From the initial state, we expand the tensor’s dimension to Eq.(2) to include backflow terms tensorbackflow . To avoid any bias introduced by UHF, optimizations of UHFs are ceased when the energy is far from fully convergence UHF .
In Fig.(1)(a)(b), red and blue solid dots denote NN and all-site backflow terms, respectively. High and low energy values for same-colored solid dots denote without and with one Lanczos optimization, respectively. Therefore, a Lanczos optimization significantly improves energy precision for either NN backflow and all-site backflow. Comparing obtained energies after one Lanczos optimization and the fPEPS with $D$=20, the NN backflow can achieve the energy precision competitive to the SU optimized fPEPS, meanwhile the all-site backflow can achieve the energy precision between SU and GO optimized fPEPS. The Tensor-Backflow has no limitations on boundary conditions. Fig.(1)(c) depicts energy results on the $4\times 16$ lattice under PBC.
To obtain the energy extrapolation under the limit of zero energy variance, we linearly fit the four solid dots for each lattice size Hubbard_TNS ; lanczos . In Fig.(1), energy extrapolations are denoted by blue solid lines. From Fig.(1)(a)(b), extrapolated energies under OBC for $8\times 8$ is -0.6993, and for $4\times 16$ is -0.6826. Extrapolated energies are close to state-of-the-art energies: -0.6993 for $8\times 8$ and -0.6830 for $4\times 16$, achieved by the fPEPS with $D$=20 within GO optimized. Based on results of OBC, the extrapolated energy of PBC in Fig.(1)(c) predicts the state-of-the-art result: -0.7642.
@@
$4\times 16$0.6797 -0.6821 -0.6830 Hubbard_fPEPS
$16\times 16$0.7219 -0.7257 -0.7260 Hubbard_fPEPS
@@
80/100 $10\times 10$1.1086 -1.1109 -1.1147 AFQMC_sym
116/144 $12\times 12$1.1038 -1.1061 -1.1102 AFQMC_sym
@@
Table 1: Energy comparisons respect to state-of-the-art results for various filling number $n$ and interaction strength $U$. A-PBC denotes the boundary condition of antiperiodic along the x-boundary and periodic along the y-boundary. $E_{p=0}$ and $E_{p=1}$ are energies achieved by Tensor-Backflow without and with one Lanczos optimization, respectively.
@@
Energy comparisons are detailed in Tab.(1). In the table, $E_{p=0}$ is from the Tensor-Backflow and $E_{p=1}$ is from a Lanczos optimization of the converged wave-function. For cases under OBC, reference energies are from GO optimized fPEPS with the largest bond dimension $D$=20 Hubbard_fPEPS . For lattices of $8\times 8$ and $4\times 16$, $E_{p=0}$ is achieved by the Tensor-Backflow with all-site backflow terms. For the lattice of $16\times 16$, $E_{p=0}$ is achieved by the Tensor-Backflow with NN backflow terms. $E_{p=1}$ are competitive to fPEPS results, and relative errors repsect to fPEPS results are $1.6\times 10^{-3}$, $1.3\times 10^{-3}$ and $4.1\times 10^{-4}$ for $8\times 8$, $4\times 16$ and $16\times 16$ lattices, respectively. Therefore, comparing to the fPEPS, energy precision for NN backflow on the $16\times 16$ lattice is even better than that for all-site backflow on 64-site lattices.
@@
Because of the high complexity, cases under PBC are not practical for fPEPS. Reference energy for $n$=0.875 and $U$=8 on the $16\times 16$ lattice under PBC is -0.7544 from the VMC method in literature Manybody_benchmark . Meanwhile, the Tensor-Backflow after a Lanczos optimization has achieved the energy $1\times 10^{-3}$ lower. Under the same filling and $U$=10, with antiperiodic along the x-boundary and periodic along the y-boundary (A-PBC in the table), the Tensor-Backflow after a Lanczos optimization has achieved the energy $2.2\times 10^{-3}$ lower comparing to the reference energy. The reference energy is from the energy extrapolation result of the tensor network method in literature Hubbard_TNS . Comparing to AFQMC results in literatures AFQMC_n1 ; AFQMC_sym , the Tensor-Backflow after a Lanczos optimization has achieved competitive energies under various fillings, interaction strengths and boundary conditions. Meanwhile, relative errors are all in the magnitude of $\mathcal{O}(10^{-3})$.
@@
Error(p=1)
NN $8\times 8$ 286720 $6.7\times 10^{-3}$ $2.3\times 10^{-3}$
$4\times 16$ 286720 $8.0\times 10^{-3}$ $3.6\times 10^{-3}$
all-site $8\times 8$ 3670016 $5.4\times 10^{-3}$ $1.6\times 10^{-3}$
$4\times 16$ 3670016 $4.9\times 10^{-3}$ $1.3\times 10^{-3}$
NN $16\times 16$ 4587520 $5.6\times 10^{-3}$ $4.5\times 10^{-4}$
@@
Table 2: Comparisons between parameter number used in Tensor-Backflow for various site number and backflow types, with $n$=0.875 and $U$=8 under OBC. Reference energies are from the GO optimized fPEPS with the bond dimension $D$=20. Parameter number is based on Eq.(2), where $Q=5,N$ for NN and all-site backflow, respectively.
Tab.(2) compares parameter number used in Tensor-Backflow for various lattice sizes and backflow types. According to Eq.(2), for the filling of $n$=0.875 on 64-site lattices, the parameter number for NN backflow and all-site backflow are 286720 and 3670016, respectively. Meanwhile, the parameter number for NN backflow on the $16\times 16$ lattice is 4587520. Although the parameter number for NN backflow on the $16\times 16$ lattice and all-site backflow on 64-site lattices are in the same magnitude, the relative error on the $16\times 16$ lattice is much lower. Because only NN hopping terms in the Fermi-Hubbard model are considered, thus the primary state representation is achieved by NN backflow terms, meanwhile the improvement of all-site backflow is moderate. For NN backflows, the parameter number grows as $\mathcal{O}(N^{2})$. When representing the Hubbard model with NN hoppings, a large parameter number of $4.6\times 10^{6}$ on the $16\times 16$ lattice has better state representation ability comparing to the small parameter number of $2.9\times 10^{5}$ on 64-site lattices.
On the $16\times 16$ lattice under OBC, the fPEPS with $D$=14 has $5.7\times 10^{6}$ parameters, and the achieved energy is -0.7224. The fPEPS with $D$=16 has $9.4\times 10^{6}$ parameters, and the achieved energy is -0.7251. The fPEPS with $D$=18 has $1.5\times 10^{7}$ parameters, and the achieved energy is -0.7260. The fPEPS with the maximal bond dimension $D$=20 has $2.2\times 10^{7}$ parameters, with the similar energy as $D$=18 Hubbard_fPEPS . For the Tensor-Backflow on the $16\times 16$ lattice, there are $E_{p=0}=$-0.7219 and $E_{p=1}=$-0.7257. Under the similar energy precision, without a Lanczos optimization, the Tensor-Backflow has many lower parameters comparing to the fPEPS with $D$=14. After a Lanczos optimization, same energy precision is achieved by the Tensor-Backflow with significantly lower parameters comparing to the fPEPS with $D$=18,20. Therefore, the Tensor-Backflow is an efficient state representation.
III.2 n=0.875
The filling of $n$=0.875 is important for understanding the mechanism of superconductivity, meanwhile the ground energies are highly competitive between various methods Hubbard_stripe .
@@
Figure 2: Patterns of ground states for $n$=0.875, $U$=8 under PBC(a)(b) and OBC(c)(d). Left-sided figures denote the average spin per site: $s_{i}^{z}=(n_{i}^{\uparrow}-n_{i}^{\downarrow})/2$, and right-sided figures denote the average density per site: $n_{i}=n_{i}^{\uparrow}+n_{i}^{\downarrow}$.
@@
On the $16\times 16$ lattice, a direct optimization for $U$=8 under PBC achieves the linear stripe order, and patterns of the ground state are depicted in Fig.(2)(a)(b). Fig.(2)(a) depicts the average spin per site: $s_{i}^{z}=(n_{i}^{\uparrow}-n_{i}^{\downarrow})/2$ and Fig.(2)(b) depicts the average density per site: $n_{i}=n_{i}^{\uparrow}+n_{i}^{\downarrow}$. As depicted by the ground state’s pattern, periods for the SDW and CDW are 16 and 8, respectively. For PBC, the ground state should be translational invariant thus an obvious pattern is not consistent with the true ground state. Since translational invariance is not enforced in the variational wave-function, even the obtained energy is competitive, the optimization is converged to a subplace of the true ground state.
For the same filling $n$ and the interaction strength $U$, we used the obtained wave-function under PBC as the initial state and then performed optimizations under OBC. The achieved state-of-the-art energy is -0.7257 as reported in Tab.(1) and (2). Meanwhile, Fig.(2)(c)(d) depict the average spin and charge density per site, respectively. Even the stripe order in the initial state has the same period as the ground state under OBC, the optimization can slightly modify the stripe’s pattern according to the boundary condition. Patterns of the ground state under OBC yield to fPEPS results Hubbard_fPEPS .
@@
Table 3: Energy comparisons between different initializations for the filling $n$=0.875. The direct optimization of $U$=8 achieves the ground state of linear stripe order, and such order is preserved when it is the initial state for other $U$ and boundary conditions. Energies with linear stripe orders are much lower than those by direct optimizations.
Figure 3: Charge and spin structure factors for the filling $n$=0.875 under various $U$ and boundary conditions. (a)(b)On the $16\times 16$ lattice, results of $U$=8 under PBC are obtained by a direct optimization, meanwhile other results are obtained by the initialization with the $U$=8 PBC wave-function. (c)(d)Results of $U$=8 under PBC by a direct optimization on the $8\times 32$ lattice.
For other interaction strengths such as $U$=10,12, Tab.(3) denote energy comparisons between the linear stripe order and direct optimizations, where linear stripe order is preserved when initializing the optimization with the wave-function obtained for $U$=8 and PBC. All energies are evaluated by $p$=1 wave-functions. In particular, direct optimization achieves diagonal stripe order when $U$=12 under PBC, and direct optimizations achieve linear SDW however no obvious CDW when $U$=10 under both PBC and A-PBC. However, the energy of linear stripe order is much lower than that of the direct optimization. Therefore, it’s concluded that the ground state is linear stripe order for large interaction strengths of $U$.
@@
$8\times 24$0.7500 -0.7546
$8\times 32$0.7505 -0.7547
$8\times 16$0.2 -0.7354 -0.7406
@@
Table 4: Energy results obtained by direct optimizations for the filling $n$=0.875 and $U$=8. $t^{\prime}$ is defined in Eq.(7). For non-zero $t^{\prime}$, NNN backflow terms are considered in the Tensor-Backflow. On the $8\times 16$ lattice, considering NNN backflow terms achieves the energy slightly lower than that when only considering NN backflow terms.
Figure 4: Patterns of ground states for $n$=0.875, $U$=8 and $t^{\prime}$=-0.2 under PBC on the $8\times 16$ lattice. Left-sided (right-sided) figure denotes the average spin (density) per site, within definitions from Fig.(2).
Figure 5: Comparisons of spin structure factors between $t^{\prime}$=0 and -0.2 for $U$=8 under PBC. Number of k-points are increased from $17^{2}$ to $65^{2}$ when calculating spin structure factors. The horizontal axis is $k_{x}$ for $t^{\prime}$=-0.2 and $k_{y}$ for $t^{\prime}$=0, meanwhile vertical axis is $S(k_{x},k_{y}=\pi)$ for $t^{\prime}$=-0.2 and $S(k_{x}=\pi,k_{y})$ for $t^{\prime}$=0.
@@
Charge and spin structure factors determine periods of the stripe order in the ground state. The structure factors are defined as: $C(\textbf{k})=\sum_{ij}e^{i\textbf{k}\cdot(\textbf{r}_{i}-\textbf{r}_{j})}n_{i}n_{j}$, $S(\textbf{k})=\sum_{ij}e^{i\textbf{k}\cdot(\textbf{r}_{i}-\textbf{r}_{j})}s_{i}^{z}s_{j}^{z}$, where $n_{i}$ and $s_{i}^{z}$ are defined in Fig.(2). $C(\textbf{k})$ is set to be zero when $k_{x}=k_{y}$=0. All results of structure factors are obtained by $p$=1 wave-functions. Fig.(3) depicts structure factors for the filling of $n$=0.875 under various $U$ and boundary conditions, on lattices of $16\times 16$ and $8\times 32$. A direct optimization for $U$=8 under PBC achieves the linear stripe order. When applies the wave-function as the initial state for $U$=10,12 under various boundary conditions, linear stripe orders are preserved. As depicted by Fig.(3)(a)(b), charge structure factor peaks at $\pi/4$ and spin structure factor peaks at $87\pi/$ 8, corresponding to periods of 8 and 16 for CDW and SDW, respectively. Fig.(3)(c)(d) depicts results of a direct optimization on the $8\times 32$ lattice for $n$=0.875 and $U$=8 under PBC. A vertical stripe order is obtained, meanwhile structure factors peaks at same locations as on the $16\times 16$ lattice. On the $16\times 16$ lattice, the lattice’s side length of 16 corresponds to a whole period of the SDW, thus peaks on the $16\times 16$ lattice are higher than those on the $8\times 32$ lattice.

### v3

@@ III.1 Optimization details
Figure 2: Energy extrapolations and comparisons respect to fPEPS Hubbard_fPEPS under OBC for the lattice of $8\times 8$(a) and $4\times 16$(b). All cases are for nn=0.875, UU=8 and $t^{\prime}$=0. Energy extrapolation for the lattice of $4\times 16$ under PBC(c) yields to the neural Pfaffian result Hubbard_NN4 . $\sigma^{2}$ is the energy variance. Red and blue solid dots depict Tensor-Backflow with NN and all-site backflow terms. For solid dots of the same color, low and high energies depict results with and without one Lanczos optimization.
III.2 Comparisons with previous results
@@
Fig.(2)(a)(b) depict energy comparisons respect to the fPEPS method under OBC on 64-site lattices. The fPEPS is a state-of-the-art method for lattice fermions and its state representation ability can be improved by increasing the bond-dimension ($D$). From literature Hubbard_fPEPS , the complexity of fPEPS scales rapidly respect to $D$: $\mathcal{O}(D^{6})$, and the maximal bond dimension is DD=20. The fPEPS is firstly optimized by the imaginary-time-evolution method named as the simple-update (SU), followed by the gradient-optimization (GO). The fPEPS with DD=20 can achieve very high energy precision under OBC.
@@
$4\times 16$ -0.6797 -0.6821 -0.6830 Hubbard_fPEPS
$16\times 16$ -0.7219 -0.7257 -0.7260 Hubbard_fPEPS
@@
80/100 $0\times 10 -1.1$086 -1.1109 -1.1147 AFQMC_sym
116/144 $2\times 12 -1.1$038 -1.1061 -1.1102 AFQMC_sym
@@
Table 1: Energy comparisons respect to state-of-the-art results for various filling number $n$ and interaction strength $U$. All cases are for $t^{\prime}$=0. A-PBC denotes the boundary condition of antiperiodic along the x-boundary and periodic along the y-boundary. $E_{p=0}$ and $E_{p=1}$ are energies achieved by Tensor-Backflow without and with one Lanczos optimization, respectively.
In our optimizations, the initial state of Tensor-Backflow is unrestricted Hartree-Fock (UHF) state represented by a tensor of the dimension $[M,N,2]$, where $M$ is the site number $N$ is the particle number and 2 is because of spin. From the initial state, we expand the tensor’s dimension to Eq.(2) to include backflow terms tensorbackflow .
In Fig.(2)(a)(b), red and blue solid dots denote NN and all-site backflow terms, respectively. High and low energy values for same-colored solid dots denote without and with one Lanczos optimization, respectively. Therefore, a Lanczos optimization significantly improves energy precision for either NN backflow and all-site backflow. Comparing obtained energies after one Lanczos optimization and the fPEPS with DD=20, the NN backflow can achieve the energy precision competitive to the SU optimized fPEPS, meanwhile the all-site backflow can achieve the energy precision competitive to GO optimized fPEPS. The Tensor-Backflow has no limitations on boundary conditions, and energies on the $4\times 16$ lattice under PBC are depicted in Fig.(2)(c).
To obtain the energy extrapolation under the limit of zero energy variance, we linearly fit the four solid dots for each lattice size Hubbard_TNS ; lanczos . In Fig.(2), energy extrapolations are denoted by blue solid lines. From Fig.(2)(a)(b), extrapolated energies under OBC for $8\times 8$ is -0.6993, and for $4\times 16$ is -0.6826. Extrapolated energies are close to state-of-the-art energies: -0.6993 for $8\times 8$ and -0.6830 for $4\times 16$, achieved by the fPEPS with DD=20 within GO optimized. Based on results of OBC, the extrapolated energy of PBC in Fig.(2)(c) is -0.7642, matching the state-of-the-art neural Pfaffian result Hubbard_NN4 .
@@
Energy comparisons respect to fPEPS Hubbard_fPEPS , NQS Manybody_benchmark , tensor networks Hubbard_TNS and AFQMC AFQMC_n1 ; AFQMC_sym are detailed in Tab.(1). In the table, $E_{p=0}$ is from the Tensor-Backflow and $E_{p=1}$ is from a Lanczos optimization of the converged wave-function. For OBC, reference energies are from GO optimized fPEPS with DD=20 Hubbard_fPEPS . On lattices of $8\times 8$ and $4\times 16$ under OBC, the wave-function’s type is all-site backflow. On the $16\times 16$ lattice under OBC, the wave-function’s type is NN backflow.
@@
In this work, the relative error of energy $E$ is defined as: $(E-E_{\mathrm{ref}})/|E_{\mathrm{ref}}|$. Under OBC, relative errors of $E_{p=1}$ respect to fPEPS results are $1.6\times 10^{-3}$, $1.3\times 10^{-3}$ and $4.1\times 10^{-4}$ for $8\times 8$, $4\times 16$ and $16\times 16$ lattices, respectively. Therefore, Tensor-Backflow has high precision for OBC. For boundary conditions other than OBC, from the VMC method in literature Manybody_benchmark , there is $E_{\mathrm{ref}}=-0.7544$ for nn=0.875 and UU=8 on the $16\times 16$ lattice under PBC, and $E_{p=1}$ is $1\times 10^{-3}$ lower. For nn=0.875 and UU=10, with antiperiodic along the x-boundary and periodic along the y-boundary, denoted by A-PBC in Tab.(1), the reference energy is from the energy extrapolation result of the tensor network method in literature Hubbard_TNS , meanwhile $E_{p=1}$ is $2.2\times 10^{-3}$ lower. In Tab.(1), for each reference energy from literatures AFQMC_n1 ; AFQMC_sym , $E_{p=1}$ is competitive and the relative error is in the magnitude of $\mathcal{O}(10^{-3})$.
$n/U$
Boundary
Condition
$t^{\prime}$
Lattice
Size
$E_{p=1}$ Ref. Hubbard_NN3
Relative
Error
0.875/8 PBC 0 $8\times 16$ -0.7568 -0.7568
$16\times 16$ -0.7552 -0.7563 $1.5\times 10^{-3}$
-0.2 $8\times 16$ -0.7410 -0.7419 $1.2\times 10^{-3}$
$12\times 12$ -0.7420 -0.7414 -$8.1\times 10^{-4}$
OBC 0 $16\times 16$ -0.7257 -0.7275 $2.5\times 10^{-3}$
Table 2: Energy comparisons respect to state-of-the-art NQS results Hubbard_NN3 for nn=0.875, UU=8 under various $t^{\prime}$, lattice sizes and boundary conditions, with NN backflow for $t^{\prime}$=0 and NNN backflow for $t^{\prime}$=-0.2. The relative error is defined as: $(E_{p=1}-E_{\mathrm{ref}})/|E_{\mathrm{ref}}|$. $E_{p=1}$ are competitive or even lower than state-of-the-art NQS results. In particular, The result for $t^{\prime}$=-0.2 on the $12\times 12$ lattice under PBC is lower than the NQS result.
From comparisons in Tab.(1), the Tensor-Backflow can achieve competitive results respect to mainstream methods. Furthremore, we compare our results to recent state-of-the-art NQS method Hubbard_NN3 in Tab.(2). In the table, except the result on OBC (detailed in Sec.III.3), each result was achieved by the first-order-gradient optimization with the step number less than $2\times 10^{4}$, without any pinning-field appied during optimizations. However, for NQS, each optimization was performed by an improved stochastic reconfiguration method with the training procedure consisting of pre-training, applying pinning-fields and the optimization step number in the magnitude of $\mathcal{O}(10^{5})$ Hubbard_NN3 . In particular, for the case of $t^{\prime}$=-0.2, nn=0.875, UU=8 on the $12\times 12$ lattice under PBC, the achieved energy is $8.1\times 10^{-4}$ lower than the NQS result, and the optimization is performed by first-order-gradient method in less than $1\times 10^{4}$ steps, optimization procedure is detailed in Sec.III.3. Therefore, the Tensor-Backflow is efficient for optimizations meanwhile has a strong state representation ability.
Tab.(3) denotes parameter numbers used in Tensor-Backflow for various lattice sizes and backflow types, as well as energy precision respect to two state-of-the-art methods under OBC. According to Eq.(2), for the filling of nn=0.875 on 64-site lattices, the parameter number for NN backflow and all-site backflow are 286720 and 3670016, respectively. Meanwhile, the parameter number for NN backflow on the $16\times 16$ lattice is 4587520. In the table, we compare $E_{p=1}$ to results achieved by state-of-the-art fPEPS Hubbard_fPEPS and NQS Hubbard_NN3 . From energy comparisons in the table, NN backflows can achieve low energy relative errors in the magnitude of $\mathcal{O}(10^{-3})$. Meanwhile, considering all backflow terms can further improve energy precision.
The parameter number of Tensor-Backflow is much lower comparing to the state-of-the-art fPEPS. On the $16\times 16$ lattice under OBC, the fPEPS with DD=14 has $5.7\times 10^{6}$ parameters, and the achieved energy is -0.7224. The fPEPS with DD=16 has $9.4\times 10^{6}$ parameters, and the achieved energy is -0.7251. The fPEPS with DD=18 has $1.5\times 10^{7}$ parameters, and the achieved energy is -0.7260. The fPEPS with the maximal bond dimension DD=20 has $2.2\times 10^{7}$ parameters, with the similar energy as DD=18 Hubbard_fPEPS . For the Tensor-Backflow on the $16\times 16$ lattice, there are $E_{p=0}=$-0.7219 and $E_{p=1}=$-0.7257. Under the similar energy precision, without a Lanczos optimization, the Tensor-Backflow has many lower parameters comparing to the fPEPS with DD=14. After a Lanczos optimization, same energy precision is achieved by the Tensor-Backflow with significantly lower parameters comparing to the fPEPS with DD=18,20. Therefore, the Tensor-Backflow is an efficient state representation.
@@
Error Hubbard_NN3
NN $8\times 8$ 286720 $2.3\times 10^{-3}$ $2.6\times 10^{-3}$
$4\times 16$ 286720 $3.6\times 10^{-3}$ $4.1\times 10^{-3}$
all-site $8\times 8$ 3670016 $1.6\times 10^{-3}$ $1.9\times 10^{-3}$
$4\times 16$ 3670016 $1.3\times 10^{-3}$ $1.8\times 10^{-3}$
NN $16\times 16$ 4587520 $4.5\times 10^{-4}$ $2.5\times 10^{-3}$
@@
Table 3: The parameter number used in Tensor-Backflow for various site number and backflow types, and comparisons of energy precision respect to state-of-the-art fPEPS Hubbard_fPEPS and NQS Hubbard_NN3 , with nn=0.875 and UU=8 under OBC. Relative errors are for $E_{p=1}$. Parameter number is based on Eq.(2), where $Q=5,M$ for NN and all-site backflows, respectively.
III.3 n=0.875
The filling of nn=0.875 is important for understanding the mechanism of superconductivity, especially when the value of $t^{\prime}$ is non-zero Hubbard_SC1 . The ground energies are highly competitive between various methods Hubbard_stripe . In this subsection, we present more details such as ground state properties and structure factors when nn=0.875 for both $t^{\prime}$=0 and $t^{\prime}$=-0.2. We firstly investigate results achieved by Tensor-Backflow when $t^{\prime}$=0.
@@
Figure 3: Patterns of ground states for nn=0.875, UU=8 and $t^{\prime}$=0 on the $16\times 16$ lattice under PBC(a)(b) and OBC(c)(d). All results are from pp=1 wave-functions with NN backflow. Left-sided figures denote the average spin per site: $s_{i}^{z}=(n_{i}^{\uparrow}-n_{i}^{\downarrow})/2$, and right-sided figures denote the average density per site: $n_{i}=n_{i}^{\uparrow}+n_{i}^{\downarrow}$.
In this work, most results from Tensor-Backflow are achieved by direct optimizations. For a direct optimization, an UHF is optimized beforehand, followed by constructing the Tensor-Backflow through expanding the tensor’s dimension to add backflow terms, meanwhile the Tensor-Backflow is optimized using either strategy from Sec.III.1. There are no pinning-fields applied during optimizations. When optimizing the UHF, the interaction strength is denoted as $U_{\mathrm{UHF}}$. Unless specifically emphasized, there is $U_{\mathrm{UHF}}=U$, and the UHF is not sufficiently optimized when $U_{\mathrm{UHF}}=U$ to avoid introducing any prior bias UHF . Futhermore, Tensor-Backflow can override properties from the initial UHF, namely, the converged wave-function is essentially a different state from the initial UHF, detailed comparisons are depicted in Appendix.(B).
@@
Fig.(3)(a)(b) depict the linear stripe order of the ground state for UU=8 on the $16\times 16$ lattice under PBC by a direct optimization. Fig.(3)(a) depicts the average spin per site: $s_{i}^{z}=(n_{i}^{\uparrow}-n_{i}^{\downarrow})/2$ and Fig.(3)(b) depicts the average density per site: $n_{i}=n_{i}^{\uparrow}+n_{i}^{\downarrow}$. As depicted by the ground state’s pattern, periods for the SDW and CDW are 16 and 8, respectively. For PBC, the ground state should be translational invariant thus an obvious pattern is not consistent with the true ground state. Since translational invariance is not enforced in the variational wave-function, even the obtained energy is competitive, the optimization is converged to a subplace of the true ground state.
@@
Table 4: Energy comparisons of $E_{p=1}$ between different initializations for nn=0.875 and $t^{\prime}$=0, with NN backflow in wave-functions. For the direct optimization of each case, the initial UHF is insufficiently optimized with $U_{\mathrm{UHF}}=U$. The direct optimization of UU=8 achieves the ground state of linear stripe order, and such order is preserved when it is transferred for other $U$ and boundary conditions. Energy convergences of transferring from UU=8 are detailed in Appendix.C. Energies with linear stripe orders are much lower than those by direct optimizations.
Figure 4: Charge and spin structure factors for nn=0.875 and $t^{\prime}$=0 under various $U$ and boundary conditions. All results are from pp=1 wave-functions, with NN backflow in wave-functions. (a)(b)On the $16\times 16$ lattice, results of UU=8 under PBC are obtained by a direct optimization, meanwhile other results are obtained by the initialization with the UU=8 PBC wave-function. (c)(d)Results of UU=8 under PBC by a direct optimization on the $8\times 32$ lattice.
For nn=0.875, UU=8, $t^{\prime}$=0 on the $16\times 16$ lattice under OBC, we transfer the wave-function converged under PBC as the initial state and then performed optimizations. The achieved state-of-the-art energy is -0.7257 as reported in Tab.(1) and (3). Meanwhile, Fig.(3)(c)(d) depict the average spin and charge density per site, respectively. Even the stripe order in the initial state has the same period as the ground state under OBC, the optimization can slightly modify the stripe’s pattern according to the boundary condition. Patterns of the ground state under OBC yield to fPEPS results Hubbard_fPEPS .
For other interaction strengths such as UU=10 and 12, Tab.(4) denote energy comparisons between the linear stripe order and direct optimizations, where linear stripe order is achieved when transferring the wave-function converged for UU=8 and PBC. In particular, direct optimization achieves diagonal stripe order when UU=12 under PBC. Ground state patterns of direct optimizations are depicted in Appendix.B. Energy convergences of transferring from UU=8 is detailed in Appendix.C. Energy comparisons are depicted in Tab.(4), and all energies are evaluated by pp=1 wave-functions. Because the energy of linear stripe order is much lower than other orders achieved by direct optimizations, the ground state is linear stripe order for large $U$.
@@
$8\times 24$ -0.7500 -0.7546
$8\times 32$ -0.7505 -0.7547
$8\times 16$ -0.2 -0.7354 -0.7406
@@
Table 5: Energy results obtained by direct optimizations for nn=0.875 and UU=8. For all cases in the table, initial UHFs are insuffieintly optimized with $U_{\mathrm{UHF}}=U$. For $t^{\prime}$=-0.2, NNN backflow terms are considered in the Tensor-Backflow. On the $8\times 16$ lattice, NNN backflow achieves the energy slightly lower than that comparing to NN backflow.
$n,U,t^{\prime}$
Backflow
Type
$U_{\mathrm{UHF}}$ $E_{p=0}$ $E_{p=1}$
0.875, 8, -0.2 NNN 0 -0.7344 -0.7395
4 -0.7370 -0.7420
8 -0.7335 -0.7388
Table 6: Comparisons of converged energies on the $12\times 12$ lattice from different $_{\mathrm{UHF}}, UU$$_{\mathrm{UHF}}$ is the interaction strength when optimizing the UHF. The UHF is fully converged when $U_{\mathrm{UHF}}\neq U$, otherwise the UHF is insufficiently optimized when $U_{\mathrm{UHF}}=U$. Initialization with $U_{\mathrm{UHF}}=4$ gives the lowest energy result for Tensor-Backflow.
@@
Structure factors determine periods of the stripe order in the ground state. The charge and spin structure factors are defined as: $C(\textbf{k})=\frac{1}{M^{2}}\sum_{ij}e^{i\textbf{k}\cdot(\textbf{r}_{i}-\textbf{r}_{j})}n_{i}n_{j}$, $S(\textbf{k})=\frac{1}{M^{2}}\sum_{ij}e^{i\textbf{k}\cdot(\textbf{r}_{i}-\textbf{r}_{j})}s_{i}^{z}s_{j}^{z}$, where $n_{i}$ and $s_{i}^{z}$ are defined in Fig.(3). $C(\textbf{k})$ is set to zero when $k_{x}=k_{y}$=0. All results of structure factors are obtained by pp=1 wave-functions. Fig.(4) depicts structure factors for the filling of nn=0.875 under various $U$ and boundary conditions, on lattices of $16\times 16$ and $8\times 32$. As depicted by Fig.(4)(a)(b), charge structure factor peaks at $\pi/4$ and spin structure factor peaks at $7\pi/8$, corresponding to periods of 8 and 16 for CDW and SDW, respectively. Fig.(4)(c)(d) depicts results of a direct optimization on the $8\times 32$ lattice for nn=0.875 and UU=8 under PBC. A vertical stripe order is obtained, meanwhile structure factors peaks at same locations as on the $16\times 16$ lattice. On the $16\times 16$ lattice, the lattice’s side length of 16 corresponds to a whole period of the SDW, thus peaks on the $16\times 16$ lattice are higher than those on the $8\times 32$ lattice.